OpenStudy (anonymous):

how many zeros does -2x^2+x-3 have

OpenStudy (anonymous):

2, it has a degree of 2, since that is its highest exponent

OpenStudy (anonymous):

how many zeros does -2x^2+x-3 have

OpenStudy (anonymous):

3x^2-9=3 solve

OpenStudy (anonymous):

3x^2-11x-4=0 factor then solve

OpenStudy (anonymous):

3x^2-11x-4=0 factor then solve

OpenStudy (anonymous):

14x^3-8x^2

OpenStudy (anonymous):

flutter y'all lol